OpenStudy (anonymous):

The length of a rectangle is 35ft, more than 4 times its width. the perimeter is 770. find its dimensions.

OpenStudy (hoblos):

l = 4w+35 2l+2w = 770 replace the first equation in the second 2(2w+35)+2w=770 8w+70 +2w = 770 10w = 700 w=70 l=4w+35 = 315